A rectangle has a width of 9 and a diagonal of 41. What is the perimeter?
Answer by edjones(8007)

a^2+b^2=c^2
a^2+9^2=41^2
a^2+81=1681
a^2=1600
a=40
2(40+9)=P
2*49=98 perimeter.
.
Ed